### v1

# Addressing the Rank Degeneration in Sequential Recommendation via Singular Spectrum Smoothing

Abstract.
Sequential recommendation (SR) investigates the dynamic user preferences modeling and generates the next-item prediction. The next-item preference is typically generated by the affinity between the sequence and item representations. However, both sequence and item representations suffer from the rank degeneration issue due to the data sparsity problem. The representations with a degenerated rank become indistinguishable due to the lacking of informative components, and thus, weaken the recommendation performance. The rank degeneration issue significantly impairs the representations for SR. This motivates us to measure how severe is the rank degeneration issue and alleviate the sequence and item representation rank degeneration issues simultaneously for SR.
In this work, we theoretically connect the sequence representation degeneration issue with the item rank degeneration, particularly for short sequences and cold items. We also identify the connection between the fast singular value decay phenomenon and the rank collapse issue in transformer sequence output and item embeddings. We propose the area under the singular value curve metric to evaluate the severity of the singular value decay phenomenon and use it as an indicator of rank degeneration. We further introduce a novel singular spectrum smoothing regularization to alleviate the rank degeneration on both sequence and item sides, which is named after the Singular sPectrum sMoothing for sequential Recommendation (SPMRec). We also establish a correlation between the ranks of sequence and item embeddings and the rank of the user-item preference prediction matrix, which can affect recommendation diversity. We conduct experiments on four benchmark datasets to demonstrate the superiority of SPMRec over the state-of-the-art recommendation methods, especially in short sequences. Further analysis reveals that our proposed method mitigates the issue of fast singular value decay, thus preserving representation rank. The experiments also demonstrate a strong connection between our proposed singular spectrum smoothing and recommendation diversity. The code is in https://github.com/zfan20/SPMRec.
Sequential Recommendation, Degeneration, Rank, Spectrum, Nuclear Norm, Singular Values
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2023; Woodstock, NYbooktitle: Recsys ’23: ACM Conference on Recommender Systems, June 03–05, 2023, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
(a) SASRec User Sequence Output Embeddings
(b) SASRec Item Embeddings
Figure 1. The normalized singular values of user sequence output and item embeddings in the four benchmark categories of Amazon Review dataset. Singular values are normalized by dividing the largest singular value.
1. Introduction
Recommender systems play a significant and widespread role in various applications, as evident in numerous scenarios. Personalized recommendations have gained immense popularity, and sequential recommendation (SR) has recently garnered considerable attention due to its scalability and superior performance. SR utilizes sequential user behaviors and item-item transition correlations to predict the user’s next preferred item. Based on the underlying sequential model architectures, SR methods consist of Markov-chain based approaches (Rendle et al., 2010; He and McAuley, 2016), convolution-based approaches (Tang and Wang, 2018; Yan et al., 2019), RNN-based approaches (Hidasi et al., 2015; Zheng et al., 2019), and Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020). Among existing SR methods, Transformer-based methods have attracted increasing interest due to their capability of modeling high-order item-item transitions and scalable architecture design.
Despite the effectiveness of existing Transformer-based methods for SR, we argue that two rank degeneration issues exist in SR: (1) sequence embeddings suffer from the rank collapse issue and (2) items have the degeneration issue in the representation learning. We conducted an empirical analysis of singular values of sequence output and item embeddings in four benchmark Amazon Review datasets by SASRec (Kang and McAuley, 2018), which is shown in Fig. (1), where fast singular value decay in both sequence and item sides is observed. We conjecture that the fast singular value decay of representations of sequence outputs and items is a principal indicator of the rank degeneration issue observed in NLP application on word embedding (Wang et al., 2020). The fast singular value decay indicates that the largest singular value dominates and other singular values are much smaller, with the potential risks of losing the embeddings rank and embeddings being distributed as a narrow cone in the latent space. The rank losing implies the lack of independent and informative components to differentiate embeddings in the latent space. We argue that the distinct long-tail distributions of sequence length and item frequency in SR cause the rank degeneration issue for both sequence and item representations learning (Gao et al., 2019; Yu et al., 2022; Qiu et al., 2022). The failure to distinguish representations undermines the recommendation performance.
We argue that the item degeneration issue introduces sequence rank collapse, implying that sequence and item degeneration issues are intertwined. Several recent works (Dong et al., 2021; Zhai et al., 2023) discuss the exponential rank collapse of the Transformer due to the self-attention module if the residual connection is removed. However, it is still an open question to clearly pinpoint the cause of the Transformer rank collapse when we consider the Transformer as a whole. We theoretically demonstrate the likelihood of sequence and item rank collapse issue in Section 4.1, which will be more severe in short sequences and cold items modeling. An intuitive understanding is that the user-item preference, calculated as the affinity between the sequence and item representations, connects the deterioration of items with the sequence degeneration. We further disclose the connection between the ranks of sequence and item representations and the diversity of the user-item preference prediction matrix, with the theoretical (Section 4.4) and empirical analyses (Section 5.8). The intuition is that the rank of the user-item preference matrix correlates with the recommendation diversity and is determined by the ranks of sequence and item embeddings. We need a proper design to regularize the sequence and item representations learning to alleviate the rank degeneration.
To this end, we propose a novel metric to measure the degree of rank degeneration issue and a novel singular spectrum smoothing regularization design to achieve better sequence and item representation learning. Specifically, as the fast singular value decay indicates the rank degeneration, we propose the Area Under the Singular value Curve (AUSC) to measure the degree of rank degeneration. Intuitively, a flat curve has a larger area under the curve than a fast decaying curve, which will also be illustrated in the following Fig. (2). We further propose the singular spectrum smoothing regularization loss as an auxiliary loss to constrain the sequence and item embedding learning, which is a surrogate of our proposed AUSC metric. We refer to the recommendation optimization with the proposed singular spectrum smoothing regularization as the Singular sPectrum sMoothing for sequential Recommendation (SPMRec). The singular spectrum smoothing regularization builds upon the nuclear norm and Frobenius norm (Horn and Johnson, 2012; Chen et al., 2022; Cui et al., 2020) to suppress the largest singular value and enlarge other singular values. Minimizing singular spectrum smoothing regularization alleviates the fast singular value decay effect and preserves the ranks of both sequence and item embeddings.
We summarize our five technical contributions as follows:
  • To the best of our knowledge, we are the first to investigate the sequence and item rank degeneration issues simultaneously in SR. We identify the connection between the fast singular value decay and the rank collapse in sequence and item embeddings and introduce a novel approach on moderating the singular value decaying phenomenon to alleviate rank collapse issues, hence avoiding embeddings being distributed in a narrow cone.
  • We propose a novel metric (AUSC) to measure the degree of rank degeneration and develop a singular spectrum smoothing regularization named SPMRec with the nuclear norm and Frobenius norm to regularize both the sequence and item embeddings.
  • We theoretically and empirically demonstrate the existence of sequence and item rank collapse issues, which are more severe in short sequences and cold items learning.
  • Our experiments on four benchmark datasets show the effectiveness of the proposed singular spectrum smoothing regularization on alleviating the fast singular value decay and preserving ranks of sequence and item embeddings.
  • We connect the rank preserving of sequence and item embeddings with the rank of the user-item preference matrix, which strongly correlates with the recommendation diversity. Our experimental results show the connection and observe that a proper contribution of smoothing improves both recommendation performance and diversity while over-emphasis deteriorates the performance significantly.
2. Related Work
2.1. Sequential Recommendation
Sequential Recommendation (SR) is a technique that encodes the temporal interactions of each user as a sequence and uses the encoded behaviors to infer their preferences. The main idea behind SR is to model the transitions between items within sequences. The earliest SR methods used Markovian approaches, such as FPMC (Rendle et al., 2010) and Fossil (He and McAuley, 2016). FPMC assumes that the item prediction depends only on the most recent previous interacted item, which is the first-order Markovian assumption. Fossil extends FPMC with the addition of item similarities information. Another line of work utilizes the Recurrent Neural Network (RNN) as the base model for modeling the temporal order of sequential interactions. The recurrent modeling characteristics of RNN make it suitable for modeling temporally ordered behaviors. Several works adopt RNNs to address the SR problem, including GRU4Rec (Hidasi et al., 2015), HGN (Ma et al., 2019; Peng et al., 2021), HRNN (Quadrana et al., 2017), and RRN (Wu et al., 2017). Different variants of RNNs are proposed, such as the original GRU in GRU4Rec and hierarchical RNN in HGN and HRNN. The third line of methods to solve the SR problem is using the Convolution Neural Network (CNN). With the interpretation of a sequence of item embeddings as an image, convolution operators are introduced to extract long-term and short-term user dynamic interests. The representative works include Caser (Tang and Wang, 2018), CosRec (Yan et al., 2019), and NextItNet (Yuan et al., 2019). Caser proposes convolutions to latent dimensions and item perspectives for comprehensive sequence encoding. Instead of treating the sequence as an image, CosRec extends to represent the sequence as a 3D tensor and adopts the 2D convolution operators. The NextItNet proposes the dilated convolution in the sequence modeling. The recent success of the self-attention module (Vaswani et al., 2017) inspires the adaptation of Transformer to the sequential recommendation task. The self-attention module captures the sequential high-order item transitions by modeling attention from all items in the sequence. SASRec (Kang and McAuley, 2018) is the earliest work of applying Transformer to SR. BERT4Rec (Sun et al., 2019) builds upon BERT (Devlin et al., 2019) with bidirectional Transformers for SR. FMLP-Rec (Zhou et al., 2022) removes the self-attention module in the Transformer and proposes an all-MLP architecture for SR. Several variants of Transformers (Fan et al., 2021; Li et al., 2020; Wu et al., 2020; Liu et al., 2021) are proposed to enhance the SR further.
2.2. Representation Degeneration
The representation degeneration problem describes the issue of learned embeddings being narrowly distributed in a small cone (Gao et al., 2019). The degeneration issue is firstly proposed in the word representations learning problem, where word embeddings are randomly initialized and then optimized via gradient descent. As shown in (Gao et al., 2019), words from text data follow the long-tail distribution, where most words have a low frequency of occurrence in the data. The long-tail distribution of words degenerates the representation learning process. Theoretically, it causes the narrow distribution of words in the representation latent space (Gao et al., 2019). Moreover, an empirical analysis (Yu et al., 2022) shows that the learning of tail words degenerates the representation optimization of frequent words. An important analysis in (Wang et al., 2020) shows that degeneration is closely related to the fast decaying singular values distribution. In other words, the largest singular value is significantly larger than all other singular values.
Distinct from the natural language processing (NLP) area, SR still has several unique characteristics. For example, sequences are short (data sparsity in recommendation systems), while sentences/paragraphs in NLP are typically long. This distinct characteristic introduces additional challenges for SR, where the degeneration issue might also happen in the sequence output embeddings. DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec and DirectAU (Wang et al., 2022) optimizes the uniformity property in the in-batch representations to enforce the representations to be uniformly distributed, where the uniformity is based on the Euclidean distance. This work connects the item embeddings degeneration with the degeneration of sequence representations. The intertwining of both sequence and item rank degeneration is rarely discussed before in SR. Moreover, some solutions from NLP to the degeneration problem only apply to the item side, including the CosReg (Gao et al., 2019) based on cosine similarities, Euclidean regularization based on learned word similarities (Zhang et al., 2020), spectrum control (Wang et al., 2020; Yan et al., 2022), and gradient gating (Yu et al., 2022). The general idea of these works is regularizing the representation matrix by strengthening the differences between representations, mostly either based on cosine similarities (e.g., CosReg (Gao et al., 2019)) or Euclidean methods (e.g., (Zhang et al., 2020) and the uniformity property adopted by DuoRec and DirectAU). Unlike existing solutions, this work investigates the rank degeneration issue from the singular value perspective of representations and proposes a novel regularization, which preserves the rank of representations and keeps representations distinguishable.
3. Preliminaries
This section presents the formal definition of the SR problem and the adoption of the Transformer in SR.
3.1. Problem Definition
In the SR problem, there is a set of users $\mathcal{U}$, a set of items $\mathcal{V}$, and interactions feedback between users and items. Each interaction is associated with its timestamp. For each user $u\in\mathcal{U}$, we can sort the interacted items chronologically and format the user as the sequence of interacted items, denoted as $\mathcal{S}^{u}={[v^{u}_{1},v^{u}_{2},\dots,v^{u}_{|\mathcal{S}^{u}|}]}$. Specifically, $v^{u}_{i}\in\mathcal{V}$ represents the $i$-th interacted item in the sequence $\mathcal{S}^{u}$ of the user $u$. SR addresses the next item prediction problem given the user history interaction sequence, which is formalized as predicting the probability of the next item $p\left(v_{|\mathcal{S}^{u}|+1}^{(u)}=v\left|\mathcal{S}^{u}\right.\right)$. In the recommendation generation stage, the next item prediction scores over all items are ranked to generate the top-N recommendation list for the user $u$.
3.2. Transformer for SR
The fundamental challenge in SR is the sequential modeling of users’ dynamic behaviors. Among existing sequential modeling methods for SR, Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019) demonstrate superior effectiveness in capturing dynamic interests and modeling different orders of item-item transitions. Following the widely adopted Transformer architecture (Vaswani et al., 2017), Transformer-based SR methods consist of several essential components, including residual connection, layer normalization, self-attention module, and feed-forward networks.
Specifically, for a given user interaction sequence $\mathcal{S}^{u}$, we pre-process it by either truncating earliest interactions or padding with zeros such that the sequence length becomes the maximum sequence length $n$ to get a fixed length sequence $s=(s_{1},s_{2},\dots,s_{n})$. An item embedding matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is randomly initialized and will be optimized, where $d$ is the latent dimension size. A trainable positional embedding $\mathbf{P}\in\mathbb{R}^{n\times d}$ is added to the sequentially ordered item embedding matrix as:
(1) $\mathbf{E}_{\mathcal{S}^{u}}=[\mathbf{m}_{s_{1}}+\mathbf{p}_{s_{1}},\mathbf{m}_{s_{2}}+\mathbf{p}_{s_{2}},\dots,\mathbf{m}_{s_{n}}+\mathbf{p}_{s_{n}}].$
Following the original Transformer (Vaswani et al., 2017), existing Transformer-based sequential methods include Feed-forward Neural Networks, Residual Connection, and Layer Normalization modules. Formally, we formulate the sequence encoder as:
(2) $\mathbf{H}_{u}=\text{SeqEnc}\left(\mathbf{E}_{\mathcal{S}^{u}}\right),$
where $\mathbf{H}_{u}\in\mathbb{R}^{n\times d}$ denotes sequential output embeddings of $\mathcal{S}^{u}$, and for each timestep $t$, and $\textbf{H}_{(u,t)}$ encodes the predicted next-item ($v_{t+1}$) representation. The next item preference score prediction is given as:
(3) $p\left(v_{t+1}=v\left|\mathcal{S}^{u}\right.\right)=\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v},$
where $\mathbf{E}_{v}$ denotes the embedding of the item $v$. The typical optimization process adopts the sampled cross-entropy loss or bayesian personalized ranking loss (BPR) (Rendle et al., 2012), which is maximizing the probability of users ranking positive items $v^{+}_{t+1}$ with greater preference scores over sampled negative items $v^{-}_{t+1}$. Specifically, the maximization of probabilistic likelihood over all users and all time steps is as follows:
(4) $\text{max}\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}p\left(v^{+}_{t+1}>v^{-}_{t+1}\left|\mathcal{S}^{u}\right.\right),$
where $v^{+}_{t+1}>v^{-}_{t+1}$ means that the positive item $v^{+}_{t+1}$ is ranked ahead of the negative item $v^{-}_{t+1}$, which can be further reformulated as a sampled cross-entropy loss minimization as:
(5) $\mathcal{L}_{\text{rec}}=\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\frac{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)}{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)+\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{-}_{t+1}}^{j}\right)},$
where $\mathbf{E}_{v^{-}_{t+1}}^{j}$ denotes the $j$-th negatively sampled item, $N^{-}_{u}$ denotes the number of negatively sampled items from the item set without interaction with the user $u$. The value of $N^{-}_{u}$ affects the recommendation performance significantly (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021). The typical choice in BPR and binary cross-entropy losses use one negatively sampled item $N^{-}_{u}=1$.
4. Singular Spectrum Smoothing
In this section, we present our proposed SPMRec framework with a novel method based on singular spectrum smoothing, with goals of alleviating the item degeneration and the Transformer sequence rank collapse issues. Specifically, we first discuss the item degeneration and Transformer rank collapse phenomenons. Then we propose a novel spectrum smoothness metric to measure the singular value curve, which indicates how severe are the degeneration and rank collapse. Then we incorporate the smoothness degree into the optimization for both item and sequence sides. We further theoretically connect our dual sides smoothing with the recommendation diversity.
4.1. Item Degeneration and Transformer Rank Collapse
As defined in Eq. (5), the next-item prediction is conducted as the dot product between the sequence output embedding $\mathbf{H}$ and the item embedding matrix $\mathbf{E}$. We argue that the item representations also have the degeneration issue, similar to the word representation degeneration issue discussed in (Gao et al., 2019; Wang et al., 2020; Zhang et al., 2020; Yu et al., 2022). Moreover, we also argue that the sequence output representations suffer from the rank collapse problem because of the degenerated item representations. The item degeneration and Transformer rank collapse show similar behavior, i.e., fast singular value decay, and we argue that item degeneration and Transformer rank collapse are closely related.
4.1.1. Item Degeneration
The item degeneration issue describes the phenomenon that most item embeddings are distributed in a narrow cone, with also the fast singular value decay of the item embedding matrix $\mathbf{E}$. As suggested in (Gao et al., 2019; Yu et al., 2022), the long-tail distribution of words in frequency is the cause of word representations degeneration. In SR, items also share the long-tail distribution. By utilizing the analysis in (Gao et al., 2019), for any tail item $v$, the recommendation loss in Eq. (5) associated with the learnable embedding $\mathbf{E}_{v}$ of the item $v$ is given as:
(6) $\min\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v}\right)+C\right),$
where $C$ includes the constant part without $\mathbf{E}_{v}$. The optimal solution is a vector with the uniformly negative direction of all sequence output embeddings $\mathbf{H}$ and infinite Frobenius norm, i.e., $||\mathbf{E}_{v}||\rightarrow\infty$, using the Theorem 1 in (Gao et al., 2019). With the long-tail distribution of item frequency, most items are cold items. Most cold items degenerate into a narrow space as a cone due to the infinite Frobenius norm. Moreover, as the Frobenius norm of item embedding matrix $||\mathbf{E}||$ is the upper bound of its largest singular value, the infinite Frobenius norm enlarges the largest singular value, causing the fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020), which is also demonstrated in the Fig. (0(b)).
4.1.2. Transformer Rank Collapse
The Transformer rank collapse problem describes the rank degeneration issue of the sequence output embedding $\mathbf{H}$. The Transformer rank degeneration issue shows the fast singular value decay phenomenon, as shown in Fig. (0(a)). Unlike the item degeneration problem, the cause of Transformer rank collapse originates from the architecture design of the Transformer. In (Dong et al., 2021), a theoretical analysis demonstrates that the self-attention module causes the rank collapse exponentially with the network depth if the residual connection is removed. A follow-up work (Zhai et al., 2023) further observes the entropy collapse in the self-attention module. We argue that the rank collapse issue still exists even with the residual connection, which is empirically shown in Fig. (0(a)).
With degenerated item embeddings, we can conclude that the short sequences output embeddings suffer from the rank degeneration. By freezing the item embedding matrix $\mathbf{E}$, we derive the rank collapse of a sequence output embedding $\mathbf{H}_{u}$ for a user $u$ in the recommendation loss as follows:
(7) $\min\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)\right)+\log\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{-}_{t+1}}^{j}\right),$
where the first component enforces the $\mathbf{H}_{u}$ to be close to the positively interacted items’ representations while the second component pulls the $\mathbf{H}_{u}$ away from negatively sampled items’ representations. When sequences are short, i.e., $|\mathcal{S}^{u}|$ is small, the second component dominates the short sequence representation learning because positively interacted items are much less than negatively sampled items. When the item degeneration happens, i.e., most items have $||\mathbf{E}_{v}||\rightarrow\infty$, the optimal solution of $\mathbf{H}_{u}$ is a vector with negative direction with most items with $||\mathbf{H}||\rightarrow 0$ as more negative items without interactions with the user $u$ are sampled with more epochs of optimization. Using the theories that the Frobenius norm is the upper bound of the largest singular value and the rank is the number of non-zero singular values, $||\mathbf{H}||\rightarrow 0$ indicates the rank losing as other singular values are being suppressed to 0.
4.2. Degree of Spectrum Smoothness
The observations in Fig. (1) show that both item embeddings and sequence output embeddings have fast singular value decay, which is a crucial indicator of the rank degeneration issue (Wang et al., 2020). It remains unknown how to measure the degree of the rank degeneration issue in the embeddings. We propose a simple, intuitive, and straightforward metric to measure the degree of spectrum smoothness, which is calculated as the Area Under Singular value Curve (AUSC). We first conduct the singular value decomposition (SVD) to decompose both sequence output embeddings $\mathbf{H}$ and item embeddings $\mathbf{E}$ as follows:
(8) $\mathbf{H}=\mathbf{U}_{\mathbf{H}}\Sigma_{\mathbf{H}}\mathbf{V}_{\mathbf{H}}^{\top},\ \ \ \mathbf{E}=\mathbf{U}_{\mathbf{E}}\Sigma_{\mathbf{E}}\mathbf{V}_{\mathbf{E}}^{\top},$
where $\mathbf{U}$ and $\mathbf{V}$ are unitary matrices, and the diagonal matrix $\Sigma$ includes singular values. We represent singular values in the $\Sigma$ as $[\sigma^{\mathbf{H}}_{1},\sigma^{\mathbf{H}}_{2},\dots,\sigma^{\mathbf{H}}_{k}]$ for sequence output embeddings $\mathbf{H}$ and $[\sigma^{\mathbf{E}}_{1},\sigma^{\mathbf{E}}_{2},\dots,\sigma^{\mathbf{E}}_{k}]$ for item embeddings $\mathbf{E}$, where singular values are in descending order as $\sigma_{1}>\sigma_{2}>\dots>\sigma_{k}>0$ and the $\sigma_{1}$ is the largest singular value $\sigma_{\text{max}}$. The area under singular curve (AUSC) is calculated as the sum of normalized singular values as follows:
(9) $\text{AUSC}=\sum_{i=1}^{k}\frac{\sigma_{i}}{\sigma_{\text{max}}}.$
When the singular value curve decays fast, as observed in the degenerated examples in Fig. (1), the area under the singular curve (AUSC) is typically small because singular values other than the largest one are all small and all singular values are normalized. We can also see that when all singular values approximate the largest singular value, i.e., $|\sigma_{i}-\sigma_{\text{max}}|\approx 0$ for all $\sigma_{i}$, the AUSC is maximized, in which the singular curve becomes flatter. For example, our toy example shown in Fig. (2) compares two singular curves, in which the left one is the fast decaying curve and the right one is an improved and more flat singular curve. The left fast decaying curve has an AUSC value of 13.64, while the right smooth curve has an AUSC value of 26.73.
Figure 2. Intuition of Spectrum Smoothing for addressing item degeneration and sequence rank collapse.
4.3. Dual Smoothing
We build upon the AUSC to propose a singular spectrum smoothing regularization as a surrogate of optimizing AUSC. With the discussion in Section 4.1, item and sequence rank degeneration issues are intertwined. Thus, we deploy the proposed singular spectrum smoothing regularization to both sequence output and item embeddings, which is dual smoothing. Specifically, the singular spectrum smoothing losses for both sequence $\mathcal{L}_{\text{seq}}$ and item $\mathcal{L}_{\text{item}}$ are conducted as follows:
(10) $\mathcal{L}_{\text{seq}}=-\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}},\ \ \ \ \ \mathcal{L}_{\text{item}}=-\frac{||\mathbf{E}||_{*}}{||\mathbf{E}||_{F}},$
where $\mathbf{H}$ is the sequence output embedding matrix of all users, and $||\mathbf{A}||_{*}$ denotes the nuclear norm as the sum of singular values for any matrix $\mathbf{A}$ ($\mathbf{A}$ can be either $\mathbf{H}$ or $\mathbf{E}$) as follows:
(11) $||\mathbf{A}||_{*}=\text{trace}\left(\left(\mathbf{A}^{\top}\mathbf{A}\right)^{\frac{1}{2}}\right)=\sum_{i}|\sigma_{i}^{\mathbf{A}}|\leq rank(\mathbf{A}),$
and $||\cdot||_{F}$ is the Frobenius norm as the upper bound of the largest singular value, $\sigma_{\text{max}}(\mathbf{A})\leq||\mathbf{A}||_{F}$ (Horn and Johnson, 2012).
Theorem 1 ().
Let $\mathbf{A}$ be a $m\times n$ matrix. Then, the nuclear norm $|A|_{*}$ is a convex relaxation of the rank($\mathbf{A}$). $|\mathbf{A}|_{*}$ provides a lower bound on the rank($\mathbf{A}$).
The singular spectrum smoothing loss is the lower bound of our proposed AUSC smoothness metric as follows:
(12) $\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}}\leq\frac{||\mathbf{H}||_{*}}{\sigma_{\text{max}}^{\mathbf{H}}},\ \ \ \ \frac{||\mathbf{E}||_{*}}{||\mathbf{E}||_{F}}\leq\frac{||\mathbf{E}||_{*}}{\sigma_{\text{max}}^{\mathbf{E}}}.$
The loss minimization of either $\mathcal{L}_{\text{seq}}$ or $\mathcal{L}_{\text{item}}$ achieves the joint effects of maximizing the nuclear norm (encouraging the sum of singular values) and minimizing the Frobenius norm (suppressing the largest singular value) and encourages more non-zero singular values, i.e., preserving the rank. With Theorem 1, the nuclear norm of $\mathbf{H}$ or $\mathbf{E}$ provides a lower bound on the rank (Horn and Johnson, 2012). When $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are minimized, the nuclear norm is maximized, improving the lower bound of the ranks of $\mathbf{H}$ and $\mathbf{E}$. We adopt the Frobenius norm instead of the largest singular value for more stable training. The minimization of $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ maximizes the AUSC of sequence embeddings and item embeddings, i.e., more smooth singular curves are introduced. Specifically, when the singular spectrum smoothing loss is minimized, the Frobenius norm is minimized so that the infinite norm of item embedding $\mathbf{E}$ is alleviated. The largest singular value is suppressed. The computational complexity of singular values calculation can be reduced when $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are conducted in batch samples, which is $O(\text{min}(B^{2}d,Bd^{2}))$ (Cui et al., 2020), and $B$ denotes the batch size and $d$ is the dimension size. Different from (Cui et al., 2021, 2020; Chen et al., 2022), we present the nuclear norm divided by the Frobenius norm on embeddings of sequence and item sides while only nuclear norm of the prediction outputs is maximized in (Cui et al., 2021, 2020; Chen et al., 2022).
4.4. Connection with Recommendation Diversity
In this section, we show the connection between our proposed dual singular spectrum smoothing and recommendation diversity. Specifically, the user-item preference matrix $\mathbf{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}|}$ is given as the dot product between sequence output embedding and item embedding matrix, which is $\mathbf{R}=\mathbf{H}^{\top}\mathbf{E}$.
The rank of $\ \mathbf{R}$ correlates with the recommendation diversity. The rank of the user-item preference prediction matrix $rank(\mathbf{R})$ indicates the number of independent components to represent the user-item preference space (Horn and Johnson, 2012; Chen et al., 2022). The rank of $\ \mathbf{R}$ equals to the rank on the user side and the rank of the item side (Horn and Johnson, 2012). From the user’s perspective, if two users have similar preferences on items, these two users are linearly dependent, and the rank reduces. This happens when two users have similar sequence output embeddings, i.e., the rank collapse of sequence output $\mathbf{H}$. From the item’s perspective, when item embedding matrix $\mathbf{E}$ degenerates, most items are narrowly distributed and hard to distinguish. The rank in the item side can be interpreted as the number of latent item topics to identify items (Koren et al., 2009).
Theorem 2 ().
[…]
With Theorem 2, we can conclude that the rank of $\mathbf{R}$ depends on the rank of sequence embedding $\mathbf{H}$ and the rank of item embeddings $\mathbf{E}$. Our proposed singular spectrum smoothing preserves ranks on both user and item sides. Therefore, the rank of the user-item preference matrix $\mathbf{R}$ is also preserved, and the recommendation diversity is correlated with our proposed smoothing method. We show the empirical relationship between item side spectrum smoothing and the recommendation diversity in Section 5.8.
4.5. Optimization with Dual Smoothing
With singular spectrum smoothing on both sequence output embeddings side $\mathcal{L}_{\text{seq}}$ and item side $\mathcal{L}_{\text{ite}}$, the final optimization loss $\mathcal{L}$ with the recommendation loss in Eq. (5) to be minimized is defined as follows:
(13) $\mathcal{L}=\mathcal{L}_{rec}+\lambda\mathcal{L}_{\text{seq}}+\beta\mathcal{L}_{\text{item}},$
where $\lambda$ is the hyper-parameter for controlling the sequence output embeddings side smoothing weight, and $\beta$ is the hyper-parameter for controlling the item side smoothing weight contribution to the final loss minimization. The final recommendation list is generated by calculating the scores defined in Eq. (3) on all items for each user. The recommendation scores on all items are sorted in descending order to produce the top-N list.
5. Experiments
We conduct several empirical studies to demonstrate the usefulness of the SPMRec and its components. We also analyze the improvements for a better understanding of the underlying mechanism. Specifically, we answer the following six research questions (RQs):
  • RQ1: Does SPMRec provide more satisfactory recommendations than state-of-the-art baselines?
  • RQ2: Does SPMRec outperform existing degeneration methods?
  • RQ3: How does each component of SPMRec contribute to the performance?
  • RQ4: Does SPMRec alleviate the rank collapse problem of Transformer?
  • RQ5: How does our item side spectrum smoothing $\beta$ affect the diversity and recommendation performance?
  • RQ6: Where do improvements of SPMRec come from?
5.1. Datasets
We present the statistics of the dataset in Table 1. We evaluate all models in four benchmarks public Amazon review datasets [^1] http://deepyeti.ucsd.edu/jianmo/amazon/index.html. In the Amazon reviews dataset, several categories of user-item interactions with timestamps are available. We choose Beauty, Toys and Games (Toys), Tools and Home (Tools), and Office Products (Office) categories in our experiments as these four categories are widely adopted benchmark categories (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022). We treat the existence of user-item reviews as positive user-item interactions. For each user, we sort the interacted items chronologically to form the user interaction sequence. The last interaction of the user sequence is used for testing and the second to last one for validation. We adopt the standard 5-core pre-processing setting on users (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022) to filter out users with less than five interactions. The detailed dataset statistics are presented in Table 1.
Table 1. Datasets Statistics.
Dataset #users #items #interactions density
avg.
interactions
per user
Beauty 22,363 12,101 198,502 0.05% 8.3
Toys 19,412 11,924 167,597 0.07% 8.6
Tools 16,638 10,217 134,476 0.08% 8.1
Office 4,905 2,420 53,258 0.44% 10.8
Total 63K 46K 554K - -
5.2. Baselines and Hyper-parameter Grid Search
We compare the proposed SPMRec with state-of-the-art sequential recommendation methods. We only include the static method BPRMF (Rendle et al., 2012) due to the page limitation. For sequential methods, we compare Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018), DT4SR (Fan et al., 2021), BERT4Rec (Sun et al., 2019), FMLP-Rec (Zhou et al., 2022), and STOSA (Fan et al., 2022). DuoRec (Qiu et al., 2022) addresses the degeneration issue in contrastive learning for SR. Note that we use only one training negative sample (i.e., $N^{-}_{u}=1$) for models with the Cross-Entropy loss for fair comparisons (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021).
We implement SPMRec with Pytorch based on the S3-Rec (Zhou et al., 2020) code base. We grid search all hyper-parameters and report the test performance based on the best validation MRR over all combinations of hyper-parameters experiments. We search the representation dimension in $\{64,128\}$, maximum sequence length from $\{50,100\}$, learning rate in $\{10^{-3},10^{-4}\}$, the $L2$ regularization control weight from $\{0.0,10^{-1},10^{-2},10^{-3}\}$, dropout rate from $\{0.3,0.5,0.7\}$ for all methods. For Transformer-based methods, we search the number of layers from $\{1,2,3\}$, and the number of heads in $\{1,2,4\}$. We adopt the early stopping strategy when the validation MRR does not increase for 50 epochs.
5.3. Evaluation
We predict the next item and generate the top-N recommendation list for each user based on the dot-product between sequence output embedding and the entire item set in descending order. We rank all items for all models so that no item sampling bias is introduced in evaluation for fair comparisons (Krichene and Rendle, 2020). The evaluation includes classical top-N ranking metrics, Recall@N and NDCG@N. Recall@N measures the percentage of sequences being recommended true preferred item in the top-N list. NDCG@N weights the ranking position of the ground truth preferred item in the ranking list by giving exponentially smaller weights to lower-ranked items. We present the averaged results over all test users, which is also the entire user set. The test results are reported based on the best validation results. We report metrics in multiple Ns, including $N=\{5,10,40\}$, which are widely adopted in most existing SR methods (Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2022).
We also evaluate the diversity of the recommendation list for all test users, which is measured as the Intra-list diversity (Zhang and Hurley, 2008; Ziegler et al., 2005) for each user’s top-N recommendation list. We focus on the top-10 ranking list for diversity calculation. The intra-list diversity is calculated as follows:
(14) $Diversity(u)=1-\frac{\sum_{i\in\mathcal{I}_{u}^{N}}\sum_{j\in\mathcal{I}_{u}^{N}}\text{cos}(\mathbf{E}_{i},\mathbf{E}_{j})}{|\mathcal{I}_{u}^{N}|*|\mathcal{I}_{u}^{N}|},$
where $\mathcal{I}_{u}^{N}$ denotes the top-$N$ recommendation list for the user $u$, $\text{cos}(\cdot,\cdot)$ represents the cosine similarity, $\mathbf{E}$ is the item embedding table defined in Section 3.2, and the reported diversity is the average diversity over all users. The larger value of diversity indicates higher intra-list diversity for the recommendation.
Table 2. Overall Performance Comparison Table. The best results are bold and the best baseline results are underlined, respectively. ‘Imp. vs. SASRec’ represents the relative improvement against SASRec without spectrum smoothing. ‘Improve.’ indicates the relative improvement against the best baseline performance.
Dataset Metric BPRMF Caser SASRec DT4SR BERT4Rec FMLP-Rec STOSA DuoRec SPMRec Imp. vs. SASRec Improv.
[…]
Toys Recall@5 0.0301 0.0240 0.0551 0.0550 $0$ul0.0625 0.0577 0.0580 0.0631 +14.52% +0.96%
[…]
5.4. Overall Comparisons (RQ1)
We present the overall performance comparison with relevant state-of-the-art baselines in Table 2. From Table 2, we can easily conclude the superiority of SPMRec in providing sequential recommendations in various datasets and metrics. We have several observations to conclude as follows:
  • SPMRec achieves the best performance and outperforms SASRec with large improvement margins. In most metrics and all datasets, SPMRec significantly outperforms the second-best baseline with 0.52% to 12.33%. The improvements demonstrate the effectiveness of SPMRec in the sequential recommendation. Furthermore, as SPMRec is built on top of the original implementation of SASRec (Kang and McAuley, 2018), we observe 6.92% to 22.53% relative improvements over SASRec with both sequence output and item embeddings spectrum smoothings. We argue that the improvements originate from the singular spectrum smoothings on both sequence and item sides.
  • SPMRec can better alleviate the sequence rank collapse issue. Comparing the AUSC of SPMRec and baselines, we can observe that SPMRec consistently achieves better AUSC values, indicating the better capability of alleviating the sequence rank issue in SR.
  • SPMRec achieves larger improvements in sparser datasets. Comparing SPMRec and SASRec, we can observe at least 17% improvements in both Beauty and Tools categories datasets, which have lower densities of 0.05% and 0.07%, respectively. However, in Toys and Office datasets with higher densities of 0.08% and 0.44%, the improvements are at most 14%. This observation demonstrates that the proposed singular spectrum smoothing SPMRec benefits sparser datasets, especially for datasets with a large portion of short sequences.
  • The consideration of Frobenius norm benefits the recommendation. STOSA adopts the sum of two L2 distances on mean and covariance embeddings and DuoRec utilizes the uniformity loss based on the exponential L2 distance. We can observe that STOSA and DuoRec both outperform the dot-product-based SASRec framework. Unlike the dot product approach, L2 distance considers the Frobenius norm of embeddings, even if it does not constrain the Frobenius norm, which again is the upper bound of the largest singular value of embeddings. This observation further demonstrates the need to consider the Frobenius norm in embedding learning. Moreover, SPMRec performs better than STOSA and DuoRec, demonstrating the necessity of considering the regularization from the singular spectrum perspective.
Table 3. Performance Comparison Table for various methods to alleviate degeneration issues.
Dataset Metric SASRec SASRec+CosReg SASRec+Euclidean SPMRec
Beauty Recall@5 0.0416 0.0410 0.0326 0.0517
NDCG@5 0.0274 0.0250 0.0212 0.0345
Recall@10 0.0633 0.0609 0.0487 0.0745
NDCG@10 0.0343 0.0314 0.0264 0.0418
Tools Recall@5 0.0284 0.0290 0.0230 0.0350
NDCG@5 0.0194 0.0194 0.0145 0.0238
Recall@10 0.0427 0.0435 0.0359 0.0513
NDCG@10 0.0240 0.0240 0.0186 0.0290
Toys Recall@5 0.0551 0.0457 0.0537 0.0631
NDCG@5 0.0377 0.0302 0.0375 0.0431
Recall@10 0.0797 0.0658 0.0717 0.0852
NDCG@10 0.0465 0.0367 0.0432 0.0503
Office Recall@5 0.0656 0.0612 0.0561 0.0714
NDCG@5 0.0428 0.0400 0.0376 0.0489
Recall@10 0.0989 0.0919 0.0862 0.1036
NDCG@10 0.0534 0.0500 0.0472 0.0593
5.5. Degeneration Methods Comparison (RQ2)
We compare the proposed SPMRec with spectrum smoothing regularization with existing methods for alleviating the degeneration issue in word embeddings learning, including the cosine regularization (Gao et al., 2019) as follows:
(15) $\mathcal{L}_{\text{item}}=\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\frac{\mathbf{E}_{i}^{\top}\mathbf{E}_{j}}{||\mathbf{E}_{i}||*||\mathbf{E}_{j}||},$
and the Euclidean method (Zhang et al., 2020) as follows:
(16) $\mathcal{L}_{\text{item}}=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{E}_{i}-\mathbf{E}_{j}||_{2}=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{E}_{i}||_{2}-2\mathbf{E}_{i}^{\top}\mathbf{E}_{j}+||\mathbf{E}_{j}||_{2}.$
Following our motivations in both sequence and item sides, we apply the cosine and Euclidean regularization methods to both sides. We report the best grid searched results of all regularization methods in Table 3. Note that the uniformity property adopted by DuoRec and DirectAU (Qiu et al., 2022; Wang et al., 2022) falls into the Euclidean methods category.
As shown in Table 3, neither cosine regularization or Euclidean regularization can consistently outperform the original SASRec without any regularization. Arguably, the potential reason is that neither of these two methods constrains the Frobenius norm of embeddings, which is the upper bound of the largest singular value, as discussed in Section 4.1. When cosine regularization and the Euclidean method are minimized, the naive solution is $||\mathbf{E}_{i}||_{2}\rightarrow\infty$ for all $i$ if we do not consider the supervised recommendation loss. Moreover, different from the word embedding task, sequences in the recommendation task are shorter on average, posing a greater challenge in addressing the degeneration and the rank collapse problem in SR. However, our proposed singular spectrum smoothing method SPMRec minimizes the Frobenius norm while maximizing the nuclear norm, achieving the goals of smoothing the singular value curve and suppressing the largest singular value simultaneously.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 3. NDCG@5 on different sequential spectrum smoothing weight $\lambda$ values. Best Viewed in colors.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 4. NDCG@5 on different item spectrum smoothing weight $\beta$ values.
5.6. Ablation Study (RQ3)
We conduct the ablation study to show the effectiveness of two important components, including singular spectrum smoothing for sequence output and item embeddings. We investigate the smoothing weights $\lambda$ (for sequence output embeddings smoothing) and $\beta$ (for item embeddings smoothing) in Fig. (3) and Fig. (4), respectively. We include the SASRec (without any regularization) and DuoRec (with uniformity regularization). Note that the special case ($\lambda=0$ and $\beta=0$) is included, which is SASRec without any regularization. Additionally, the sequence outputs and item embeddings spectrum smoothing losses are in different scales ($10^{-1}$ for sequence embeddings and $10^{-4}$ for item embeddings). Hence, the scales of $\lambda$ and $\beta$ are different with the purpose of balancing the contributions of two regularizations.
5.6.1. Sequence Output Embedding Smoothing Weight $\lambda$
We present the performance values with different values of $\lambda$ in Fig. (3) for all four datasets. Fig. (3) shows that the proper selection of $\lambda$ improves the performance significantly. Moreover, most $\lambda$ values achieve comparative and better performances than SASRec. These observations show that the sequence output embedding spectrum smoothing is beneficial for the sequential recommendation.
5.6.2. Item Embedding Smoothing Weight $\beta$
The NDCG@5 values of different $\beta$ are shown in Fig. (4), which is controlling the spectrum smoothing component’s contribution. The analysis is conducted in all four datasets used. Overall, when $\beta$ becomes larger, i.e., having a larger contribution to the overall loss, the performance increases. It shows the effectiveness of item embeddings singular spectrum smoothing for performance improvements.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 5. Normalized Singular Values Curve Comparison between SASRec and the proposed SPMRec in best grid searched settings.
5.7. Sequential User Singular Values Decay Analysis (RQ4)
We visualize the singular values curve of sequence output embedding for our proposed SPMRec and SASRec, as shown in Fig. (5). We can observe that the proposed SPMRec has a smoother decaying curve than SASRec. Moreover, for SPMRec, the degrees of spectrum smoothness for all four datasets are 26.34 for Beauty, 20.76 for Tools, 20.25 for Toys, and 19.24 for Office. However, SASRec has lower degrees for spectrum smoothness as 19.54 for Beauty, 17.10 for Tools, 17.00 for Toys, and 13.67 for Office. SPMRec consistently has larger degrees of spectrum smoothness than SASRec, which demonstrates the effectiveness of SPMRec in preventing Transformer from rank collapse for the sequential recommendation.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 6. Relationship between item spectrum smoothing weight $\beta$ and recommendation diversity. Best Viewed in colors.
5.8. Diversity and Item Spectrum Smoothing (RQ5)
Item Spectrum Smoothing controlled by $\beta$ is closely related to the recommendation diversity. Moreover, the recommendation performance drops when we demand a high recommendation diversity. We demonstrate a strong connection between our item spectrum smoothing and the recommendation diversity by showing the relationship of $\beta$ and the diversity, as shown in Fig. (6). In Fig. (6), the x-axis denotes the value of $\beta$. As the $\beta$ becomes larger, we can observe that the diversity consistently increases, in all four datasets, demonstrating the strong connection between the recommendation diversity and our proposed item spectrum smoothing component. Another observation is that the NDCG@10 drops significantly when we enforce too large $\beta$, which also indicates a higher diversity. This observation demonstrates that a proper selection of $\beta$ can achieve satisfactory recommendation performance and diversity simultaneously. However, demanding higher diversity can hurt the recommendation performance.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 7. NDCG@5 on different sequence groups based on sequence length.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 8. NDCG@5 on different item groups based on item popularity.
5.9. Improvements Analysis (RQ6)
In this section, we investigate the improvements of SPMRec from both sequence and item sides for a better understanding of the advantages of SPMRec. The improvements analysis are shown in Fig. (7) and Fig. (8), for sequences and items respectively. We compare with three strong baselines, including SASRec, STOSA, and DuoRec (having data augmentations).
5.9.1. Improvements w.r.t. Sequence Length
SPMRec is highly effective in short sequences recommendation. We group sequences (users) based on the number of training interactions (sequence length) of users. Within each group, we average the testing performances of sequences in the group. The comparisons on each group are shown in Fig. (7). We can observe that SPMRec achieves significant improvements in short sequences. Specifically, the improvements over the best STOSA range from 7.21% to 23.04%. The recommendation for short sequences is challenging due to limited interaction signals, which is known as the cold start problem. SPMRec regularizes the sequence output embeddings and prevents the rank collapse so that short sequences can be distinguishable and are properly modeled. We can also observe that SPMRec has inferior performances in the longest sequence group, which is even worse than SASRec in the Toys dataset. However, the short sequences are in majority in both the used datasets and practical applications. With the proposed singular spectrum smoothing, SPMRec benefits short sequences with the sacrifices of long sequences.
5.9.2. Improvements w.r.t. Item Popularity
The SPMRec focuses the representation learning for popular items with the best searched $\beta$. We group items based on the number of training reviews of items, i.e., item popularity. We report the average testing performance of items in each group. We show the comparisons in Fig. (8). We can observe that SPMRec achieves significant improvements in the most popular groups of items while only comparative and even inferior performances in less popular groups. Fig. (8) shows the group performance with the best searched $\beta$, which emphasizes popular items. However, this observation does not contradict the motivation of item embedding singular spectrum smoothing as we observe that a large $\beta$ provides more balanced performances in different item groups while inferior overall sequences recommendation performances.
6. Conclusion
We investigate the rank degeneration issue in both sequence output embeddings and item embeddings, which are observed as the fast singular value decay phenomenon. We argue that the rank degeneration issues in both sequences and items are intertwined. With theoretical demonstrations, we show the possibility of sequence rank collapse in SR, which is more severe in short sequence learning. We propose a novel metric to measure the degree of rank degeneration, which is calculated as the area under the singular value curve (AUSC). We further propose a singular spectrum smoothing regularization as a surrogate of AUSC to constrain the Transformer and item embeddings learning, alleviating the rank collapse issue and preserving the ranks of embeddings. We also connect the ranks of sequence and item embeddings with the rank of the user-item preference prediction matrix, which correlates with the recommendation diversity. We empirically demonstrate the effectiveness of the proposed singular spectrum smoothing regularization in recommendation performances, especially for short sequences and popular items. We also observe a strong connection between the proposed singular spectrum smoothing regularization and the recommendation diversity.