### v2

# Sequential Recommendation with Controllable Diversification: Representation Degeneration and Diversity

Abstract.
Sequential recommendation (SR) models the dynamic user preferences and generates the next-item prediction as the affinity between the sequence and items, in a joint latent space with low dimensions (i.e., the sequence and item embedding space). Both sequence and item representations suffer from the representation degeneration issue due to the user/item long-tail distributions, where tail users/items are indistinguishably distributed as a narrow cone in the latent space. We argue that the representation degeneration issue is the root cause of insufficient recommendation diversity in existing SR methods, impairing the user potential exploration and further worsening the echo chamber issue.
In this work, we first disclose the connection between the representation degeneration and recommendation diversity, in which severer representation degeneration indicates lower recommendation diversity. We then propose a novel Singular sPectrum sMoothing regularization for Recommendation (SPMRec), which acts as a controllable surrogate to alleviate the degeneration and achieve the balance between recommendation diversity and performance. The proposed smoothing regularization alleviates the degeneration by maximizing the area under the singular value curve, which is also the diversity surrogate. We conduct experiments on four benchmark datasets to demonstrate the superiority of SPMRec, and show that the proposed singular spectrum smoothing can control the balance of recommendation performance and diversity simultaneously.
Recommendation, Nuclear Norm, Determinant, Diversity
(a) Beauty
(b) Toys
Figure 1. A demonstrated strong correlation between the proposed spectrum smoothing regularization (weighted by $\beta$) and intra-list diversity on public Amazon review data. A challenging balance between recommendation performance and diversity can be achieved with the proposed method.
1. Introduction
Recommender systems play a significant and widespread role in providing personalized recommendations and user modeling for various web applications, as evident in fashion, music, movies, and marketing scenarios. Among existing frameworks, sequential recommendation (SR) (Rendle et al., 2010; He and McAuley, 2016; Tang and Wang, 2018; Hidasi et al., 2015; Kang and McAuley, 2018) has recently garnered considerable attention due to its scalability and superior performance. SR utilizes sequential user behaviors and item-item transition correlations to predict the user’s next preferred item. Among existing SR methods, Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) have attracted increasing interests due to their capability of modeling high-order item-item transitions and scalable architecture design. In addition to recommendation quality (quantified as next-item prediction accuracy), a diverse set of items generation is also crucial for improving user satisfaction (Zhang and Hurley, 2008; Chen et al., 2018) and helping users escape the echo chamber (Ge et al., 2020) for better long-term engagements (Chen et al., 2021).
The long-tail distribution of users/items has been widely observed in real-world applications and benchmarking datasets (Zhang et al., 2023; Domingues et al., 2012; Wang et al., 2022a), in which a small portion of users/items contribute to most interactions. According to empirical observations (Goel et al., 2010), tail items availability in recommendation can boost the sales of head items, which triggers the need of diverse recommendations. Demonstrated by (Wang et al., 2018; Gao et al., 2023), the long-tail distribution potentially leads to more intense Matthew effect, in which popular items and categories further dominate the recommender system with the implication of insufficient recommendation diversity. However, few research investigates the relationship between the long-tail distribution of users/items and the recommendation diversity.
Existing methods either focus on post re-ranking approaches (with the risk of sacrificing recommendation quality) (Zhang and Hurley, 2008; Chen et al., 2018; Wilhelm et al., 2018) or build upon the indirect connection between recommendation quality and diversity (Yang et al., 2023; Liu et al., 2022). A balance between recommendation quality and diversity is needed because over-emphasizing either the quality or the diversity leads to sub-optimal user satisfaction. To achieve this balance, a recommendation model with controllable diversification module is needed. It is also challenging to develop the controllable diversity module because the underlying connection between the quality and the diversity is not identified yet, and a differentiable solution is preferred so as to optimize both the quality and diversity.
(a) User Sequence Embeddings
(b) Item Embeddings
Figure 2. Fast singular values decay of user sequence output and item embeddings of SASRec are observed, implying the representation degeneration exists in both users and items. Singular values are normalized by dividing the largest singular value.
To this end, we demonstrate the connection between the long-tail distribution on users/items and the recommendation diversity, via the observed representation degeneration issue in both users and items, as shown in Fig. (2). We theoretically show that the fast singular value decay from the users/items representation degeneration is the root cause of insufficient diversity. More details will be explained in Section  3. Motivated by this connection, we develop a novel sequential recommendation with controllable diversity module by smoothing the singular value curves for users/items representations, with the goals of alleviating the degeneration and improving the recommendation diversity simultaneously.
Specifically, we propose a novel metric to measure the degree of degeneration issue and a novel Singular sPectrum sMoothing regularization design to achieve better sequence and item representation learning for Recommendations, SPMRec. We propose the Area Under the Singular value Curve (AUSC) to measure the degree of degeneration, with the intution that a flat curve has a larger area under the curve than a fast decaying curve, which will also be illustrated in the following Fig. (3). We further propose the singular spectrum smoothing regularization loss as an auxiliary loss to constrain the sequence and item embedding learning, which is a surrogate of our proposed AUSC metric. The singular spectrum smoothing regularization builds upon the nuclear norm and Frobenius norm (Horn and Johnson, 2012; Chen et al., 2022; Cui et al., 2020) to suppress the largest singular value and enlarge other singular values. The singular spectrum smoothing regularization unblocks the feasibility of sequential recommendation with controllable diversity. We summarize our three technical contributions as follows:
  • To the best of our knowledge, we are the first to identify the connection between the representation degeneration and insufficient diversity in SR from the theoretical perspective.
  • We propose a novel metric (AUSC) to measure the degree of degeneration and develop a singular spectrum smoothing regularization named SPMRec to achieve the important but challenging balance of recommendation quality and diversity.
  • Our experiments on four benchmark datasets show the effectiveness of the proposed SPMRec, and empirically demonstrate that a proper contribution of smoothing improves both recommendation performance and diversity.
2. Preliminaries
2.1. Problem Definition
In the SR problem, there are a set of users $\mathcal{U}$, a set of items $\mathcal{V}$, and interactions feedback between users and items. Each interaction is associated with its timestamp. For each user $u\in\mathcal{U}$, we sort the interacted items chronologically and format the user as the sequence of interacted items, denoted as $\mathcal{S}^{u}={[v^{u}_{1},v^{u}_{2},\dots,v^{u}_{|\mathcal{S}^{u}|}]}$. Specifically, $v^{u}_{i}\in\mathcal{V}$ represents the $i$-th interacted item in the sequence $\mathcal{S}^{u}$ of the user $u$. SR tackles the problem of predicting the next item based on the sequence of user’s historical interactions, which is formalized as predicting the probability of the next item $p\left(v_{|\mathcal{S}^{u}|+1}^{(u)}=v\left|\mathcal{S}^{u}\right.\right)$. In the recommendation generation stage, the next item prediction scores over all items are ranked to generate the top-N recommendation list for the user $u$.
2.2. Sequential User Modeling
Given a user interaction sequence $\mathcal{S}^{u}$, we pre-process it by either truncating earliest interactions or padding with zeros such that the sequence length becomes the maximum sequence length $n$ to get a fixed length sequence $s=(s_{1},s_{2},\dots,s_{n})$. An item embedding matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is randomly initialized and will be optimized, where $d$ is the latent dimension size. A trainable positional embedding $\mathbf{P}\in\mathbb{R}^{n\times d}$ is added to the sequentially ordered item embedding matrix as $\mathbf{E}_{\mathcal{S}^{u}}$. Formally, we formulate the sequence encoder as follows:
(1) $\mathbf{H}_{u}=\text{SeqEnc}\left(\mathbf{E}_{\mathcal{S}^{u}}\right),$
where $\mathbf{H}_{u}\in\mathbb{R}^{n\times d}$ denotes sequential output embeddings of $\mathcal{S}^{u}$, and for each timestep $t$, and $\textbf{H}_{(u,t)}$ encodes the predicted next-item ($v_{t+1}$) representation.
2.3. Next-Item Prediction Optimization
The next item preference score prediction is given as:
(2) $p\left(v_{t+1}=v\left|\mathcal{S}^{u}\right.\right)=\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v},$
where $\mathbf{E}_{v}$ denotes the embedding of the item $v$. The typical optimization process adopts the classical sampled cross-entropy (sampled CE) loss or bayesian personalized ranking loss (BPR) (Rendle et al., 2012), which is maximizing the probability of users ranking positive items $v^{+}_{t+1}$ with greater preference scores over sampled negative items $v^{-}_{t+1}$. The optimization can be formulated as a sampled cross-entropy loss minimization as:
(3) $\mathcal{L}_{\text{rec}}=\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\frac{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{+}_{t+1}}\right)}{\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{-}_{t+1}}^{j}\right)},$
where $\mathbf{M}_{v^{-}_{t+1}}^{j}$ denotes the $j$-th negatively sampled item, $N^{-}_{u}$ denotes the number of negatively sampled items from the item set without interaction with the user $u$. The value of $N^{-}_{u}$ affects the recommendation performance significantly (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021). The typical choice in BPR and binary cross-entropy losses uses one negatively sampled item $N^{-}_{u}=1$.
3. Representation Degeneration and Diversity
3.1. Representation Degeneration Revisit
3.1.1. Revisit of Item Degeneration
Most item embeddings are distributed in a narrow cone when the item degeneration issue occurs, with observation as fast singular value decay of the item embedding matrix $\mathbf{M}$ (Yu et al., 2022; Qiu et al., 2022). As suggested in (Gao et al., 2019; Yu et al., 2022), the long-tail distribution in frequency is the cause of representations degeneration. In SR, items follow the long-tail distribution. By utilizing the analysis in (Gao et al., 2019), for any tail item $v$, the recommendation loss in Eq. (3) associated with the learnable embedding $\mathbf{M}_{v}$ of the item $v$ is given as:
(4) $\min\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v}\right)+C\right),$
where $C$ includes the constant part without $\mathbf{M}_{v}$. The optimal solution is a vector with the uniformly negative direction of all sequence output embeddings $\mathbf{H}$ and infinite Frobenius norm, i.e., $||\mathbf{M}_{v}||\approx\infty$, using the Theorem 1 in (Gao et al., 2019). With the long-tail distribution of item frequency, most items are cold items. Most cold items degenerate into a narrow space as a cone, with the observation of fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020) and Fig. (2(b)). Moreover, as the Frobenius norm of item embedding matrix $||\mathbf{E}||$ is the upper bound of its largest singular value, the infinite Frobenius norm enlarges the largest singular value, causing the fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020), which is also demonstrated in the Fig. (2(b)).
3.1.2. Item Degeneration Induces Sequence Degeneration
With degenerated item embeddings, we conclude that output embeddings of short sequences suffer from the degeneration issue. By freezing the item embedding matrix $\mathbf{M}$, we derive the recommendation loss of a sequence output embedding $\mathbf{H}_{u}$ for user $u$ as follows:
(5) $\min\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{+}_{t+1}}\right)\right)+\log\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{-}_{t+1}}^{j}\right),$
where the first component enforces the $\mathbf{H}_{u}$ to be close to the positively interacted items’ representations while the second part pulls the $\mathbf{H}_{u}$ away from negatively sampled items’ representations.
When the item degeneration happens, i.e., most tail items have $||\mathbf{M}_{v}||\approx\infty$, the optimal solution of $\mathbf{H}_{u}$ is a vector in the negative direction with most items, indicating $||\mathbf{H}_{u}||\approx 0$. This is because more negative items are sampled with more optimization steps. $||\mathbf{H}_{u}||\rightarrow 0$ indicates other singular values are being suppressed to 0, based on theories that the Frobenius norm is the upper bound of the largest singular value. When $|\mathcal{S}^{u}|\approx 0$ (i.e., sequences are short), the degeneration issue becomes more severe, as there are much more negatively sampled items than positive items, where the second component dominates the learning process of $\mathbf{H}_{u}$.
3.2. Theoretical Connection with Diversity
3.2.1. Determinant of Embedding as Diversity Measurement
We connect the determinant of embedding with the diversity via the well-defined determinant maximization problem in the determinant point process (DPP) (Kulesza et al., 2012). The determinant maximization problem (Kulesza et al., 2012) has been studied extensively in several machine learning applications, such as core-set discovery (Mahabadi et al., 2019; Mirrokni and Zadimoghaddam, 2015), feature selection (Zadeh et al., 2017), and recommendation re-ranking (Wilhelm et al., 2018; Liu et al., 2022; Chen et al., 2018; Huang et al., 2021). Specifically, the determinant maximization problem is to find a subset from the total item set $\mathcal{Y}\in\mathcal{V}$ with $|\mathcal{Y}|\ll|\mathcal{V}|$. Formally, given the item feature matrix $\mathbf{F}\in\mathbb{R}^{|\mathcal{V}|\times d}$, DPP finds an item subset $\mathcal{Y}$, such that the determinant of the kernel matrix of the item subset $\text{det}(\mathbf{K}_{\mathcal{Y}})$ is maximized as follows:
(6) $\operatorname*{arg\,max}_{\mathcal{Y}\in\mathcal{V}}\text{det}(\mathbf{K}_{\mathcal{Y}}),$
where the kernel $\mathbf{K}_{\mathcal{Y}}(\cdot,\cdot)\in\mathbb{R}^{|\mathcal{Y}|\times|\mathcal{Y}|}$ is typically defined as the dot product of the item features $\mathbf{K}_{\mathcal{Y}}(\mathcal{Y},\mathcal{Y})=\mathbf{F}_{\mathcal{Y}}^{\top}\mathbf{F}_{\mathcal{Y}}$. The existing DPP solutions (Mahabadi et al., 2019; Mirrokni and Zadimoghaddam, 2015; Wilhelm et al., 2018) iteratively expand the diverse subset by searching the item to enlarge the determinant of kernel $\text{det}(\mathbf{K}_{\mathcal{Y}})$.
A subset with two items case (i.e., $\mathcal{Y}=\{v_{i},v_{j}\}$) can be used to demonstrate the relationship between determinant of the kernel $\mathbf{K}_{\mathcal{Y}}$ and the diversity of the subset $\mathcal{Y}$. The determinant of the kernel $\text{det}(\mathbf{K}_{\mathcal{Y}})$ can be extended as $\text{det}(\mathbf{K}_{\mathcal{Y}})=\mathbf{K}_{v_{i}v_{i}}\mathbf{K}_{v_{j}v_{j}}-\mathbf{K}_{v_{i}v_{j}}\mathbf{K}_{v_{j}v_{i}}$, where $\mathbf{K}_{v_{i}v_{j}}\mathbf{K}_{v_{j}v_{i}}$ identifies the correlations with different items. When $\text{det}(\mathbf{K}_{\mathcal{Y}})$ obtains the minimum value 0, $\mathbf{K}_{v_{i}v_{j}}=\mathbf{K}_{v_{i}v_{i}}=\mathbf{K}_{v_{j}v_{j}}$, which indicates that $v_{i}$ and $v_{j}$ are highly correlated, i.e., no diversity. On the contrary, when $\text{det}(\mathbf{K}_{\mathcal{Y}})$ is maximized, the diverse set of items is retrieved.
Extending to the total item set $\mathcal{V}$ for recommendations in SR, the total item set’s diversity can be measured as the determinant of the kernel on $\mathcal{V}$ and the associated item embedding matrix $\mathbf{M}$, which is $\text{det}(\mathbf{K}_{\mathcal{V}})=\text{det}(\mathbf{M}^{\top}\mathbf{M})$. Inductively, given a selected item set with the size as $|\mathcal{Y}|=n$, DPP adds an item $v_{j}$ to the selected item set $\mathcal{Y}$ as $\hat{\mathcal{Y}}$ with size $n+1$. Using Matrix Determinant Lemma, the the determinant of $\hat{\mathcal{Y}}$ is given as:
(7) $\text{det}(\hat{\mathcal{Y}})$ $=\text{det}\left(\begin{bmatrix}\mathbf{K}_{\mathcal{Y}}&\mathbf{K}_{\mathcal{Y},v_{j}}=\mathbf{F}_{\mathcal{Y}}^{\top}\mathbf{F}_{v_{j}}\\ \mathbf{K}_{v_{j},\mathcal{Y}}=\mathbf{F}_{v_{j}}^{\top}\mathbf{F}_{\mathcal{Y}}&\mathbf{K}_{v_{j},v_{j}}=\mathbf{F}_{v_{j}}^{\top}\mathbf{F}_{v_{j}}\end{bmatrix}\right)$
(8) $=\text{det}(\mathbf{K}_{\mathcal{Y}})\cdot\text{det}\left(\mathbf{K}_{v_{j},v_{j}}-\frac{\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}}{\mathbf{K}_{\mathcal{Y}}}\right).$
$\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}$ measures the correlations between the item $v_{j}$ and the selected item set $\mathcal{Y}$. As $\mathcal{Y}$ is selected and fixed, $\text{det}(\mathbf{K}_{\mathcal{Y}})$ and $\mathbf{K}_{\mathcal{Y}}$ can be treated as constants, and $\mathbf{K}_{v_{j},v_{j}}$ is the similarity of the item itself and can also be treated as a constant. DPP aims at finding the $v_{j}$ for maximizing the $\text{det}(\hat{\mathcal{Y}})$, which is equivalent to finding the $v_{j}$ with the least correlations with the selected item set $\mathcal{Y}$:
(9) $\operatorname*{arg\,max}_{v_{j}}\text{det}(\mathbf{K}_{\hat{\mathcal{Y}}})=\operatorname*{arg\,min}_{v_{j}}\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}.$
This inductive argument implies that DPPs favor a diverse item set.
By applying the logorithmic and the SVD decomposition on the embedding matrix $\mathbf{M}$, we obtain the following the relationship between diversity and the singular values of $\mathbf{M}$:
(10) $\log\text{det}(\mathbf{M}^{\top}\mathbf{M})=\log\prod_{i=1}^{d}\left(\sigma^{\mathbf{M}}_{i}\right)^{2}=\sum_{i=1}^{d}2\log\sigma^{\mathbf{M}}_{i}\propto\sum_{i=1}^{d}\sigma^{\mathbf{M}}_{i},$
where $\left[\sigma^{\mathbf{M}}_{1},\sigma^{\mathbf{M}}_{2},\dots,\sigma^{\mathbf{M}}_{d}\right]$ are singular values of the item embedding matrix $\mathbf{M}$, and singular values are in the descending order as $\sigma^{\mathbf{M}}_{1}>\sigma^{\mathbf{M}}_{2}>\dots>\sigma^{\mathbf{M}}_{d}>0$ and the $\sigma^{\mathbf{M}}_{1}$ is the largest singular value $\sigma^{\mathbf{M}}_{\text{max}}$. With Eq. (10), we establish the theoretical relationship between the determinant of embedding and item set diversity. As the 2log only re-scales the singular value, we have shown that the diversity is proportional to the sum of singular values, i.e., the nuclear norm of the embedding $\sum_{i=1}^{d}\sigma^{\mathbf{M}}_{i}=|\mathbf{M}|_{*}$.
Figure 3. Intuition of Singular Spectrum Smoothing for addressing sequence and item representation degeneration.
3.2.2. Representation Degeneration and Insufficient Diversity
The observations in Fig. (2) show that both item embeddings and sequence output embeddings find the fast singular value decay, which is a crucial indicator of the representation degeneration issue (Wang et al., 2020). With SVD to decompose both sequence output embeddings $\mathbf{H}$ and item embeddings $\mathbf{M}$, we obtain singular values of $\mathbf{H}$ and $\mathbf{M}$ as $\left[\sigma^{\mathbf{H}}_{1},\sigma^{\mathbf{H}}_{2},\dots,\sigma^{\mathbf{H}}_{d}\right]$ and $\left[\sigma^{\mathbf{M}}_{1},\sigma^{\mathbf{M}}_{2},\dots,\sigma^{\mathbf{M}}_{d}\right]$.
The fast singular value decay implies that the first singular value is significantly larger than others, i.e., $\sigma^{\mathbf{H}}_{1}\gg\sigma^{\mathbf{H}}_{2}$ and $\sigma^{\mathbf{M}}_{1}\gg\sigma^{\mathbf{M}}_{2}$. It further implies the insufficient diversity of sequence and item embeddings because the determinant (re-scaled sum of singular values) of the embedding, i.e., the diversity from the Eq. (10), achieves a small value when the fast singular value decay happens (most singular values are small values).
4. Singular Spectrum Smoothing
Motivated by the theoretical connection between the representation degeneration and diversity via the determinant of embedding, we present our proposed SPMRec framework with a novel metric to measure the degree of representation degeneration and a novel singular spectrum smoothing regularization for alleviating the degeneration issue and controlling the diversity simultaneously.
4.1. Degree of Spectrum Smoothness
We propose a simple, intuitive, and straightforward metric to measure the degree of singular spectrum smoothness, which is calculated as the Area Under Singular value Curve (AUSC). The area under singular curve (AUSC) is calculated as the sum of normalized singular values as follows:
(11) $\text{AUSC}=\sum_{i=1}^{k}\frac{\sigma_{i}}{\sigma_{\text{max}}}.$
When the singular value curve decays fast, as observed in the degenerated examples in Fig. (2), the area under the singular curve (AUSC) is typically small because singular values other than the largest one are all small and all singular values are normalized. We can also see that when all singular values approximate the largest singular value, i.e., $|\sigma_{i}-\sigma_{\text{max}}|\approx 0$ for all $\sigma_{i}$, the AUSC is maximized, in which the singular curve becomes flatter. For example, our toy example shown in Fig. (3) compares two singular curves, in which the left one is the fast decaying curve and the right one is an improved and more flat singular curve. The left fast decaying curve has an AUSC value of 13.64, while the right smoothed curve has an AUSC value of 26.73.
4.2. Dual Smoothing
We build upon the AUSC to propose a singular spectrum smoothing regularization as a surrogate of optimizing AUSC. With the discussion in Section 3.1, item and sequence degeneration issues are intertwined. Thus, we deploy the proposed singular spectrum smoothing regularization to both sequence output and item embeddings, which is referred to as dual smoothing. Specifically, the singular spectrum smoothing losses for both sequence $\mathcal{L}_{\text{seq}}$ and item $\mathcal{L}_{\text{item}}$ are conducted as follows:
(12) $\mathcal{L}_{\text{seq}}=-\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}},\ \ \ \ \ \mathcal{L}_{\text{item}}=-\frac{||\mathbf{M}||_{*}}{||\mathbf{M}||_{F}},$
where $\mathbf{H}$ is the sequence output embedding matrix of all users, and The singular spectrum smoothing loss is the lower bound of our proposed AUSC smoothness metric as follows:
(13) $\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}}\leq\frac{||\mathbf{H}||_{*}}{\sigma_{\text{max}}^{\mathbf{H}}},\ \ \ \ \frac{||\mathbf{M}||_{*}}{||\mathbf{M}||_{F}}\leq\frac{||\mathbf{M}||_{*}}{\sigma_{\text{max}}^{\mathbf{M}}}.$
The loss minimization of either $\mathcal{L}_{\text{seq}}$ or $\mathcal{L}_{\text{item}}$ achieves the joint effects of
  • maximizing the nuclear norm (encouraging the sum of singular values),
  • minimizing the Frobenius norm (suppressing the largest singular value),
  • and enlarges more tail singular values.
When $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are minimized, the nuclear norm is maximized, improving the determinant and diversity of $\mathbf{H}$ and $\mathbf{M}$. We adopt the Frobenius norm instead of the largest singular value for more stable training. The minimization of $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ maximizes the AUSC of sequence embeddings and item embeddings, i.e., more smooth singular curves are introduced. Specifically, when the singular spectrum smoothing loss is minimized, the Frobenius norm is minimized so that the infinite norm of item embedding $\mathbf{M}$ is alleviated. The largest singular value is suppressed.
Computational Efficiency: the computational complexity of singular values calculation can be reduced when $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are conducted in batch, which is $O(\text{min}(B^{2}d,Bd^{2}))$ (Cui et al., 2020), and $B$ denotes the batch size and $d$ is the dimension size. Different from (Cui et al., 2021, 2020; Chen et al., 2022), we present the nuclear norm divided by the Frobenius norm on embeddings of sequence and item sides while only nuclear norm of the prediction outputs is maximized in (Cui et al., 2021, 2020; Chen et al., 2022).
4.3. Optimization with Dual Smoothing
With singular spectrum smoothing on both sequence output embeddings side $\mathcal{L}_{\text{seq}}$ and item side $\mathcal{L}_{\text{item}}$, the final optimization loss $\mathcal{L}$ with the recommendation loss in Eq. (3) to be minimized is defined as follows:
(14) $\mathcal{L}=\mathcal{L}_{rec}+\lambda\mathcal{L}_{\text{seq}}+\beta\mathcal{L}_{\text{item}},$
where $\lambda$ is the hyper-parameter for controlling the sequence output embeddings side smoothing weight, and $\beta$ is the hyper-parameter for controlling the item side smoothing weight contribution to the final loss minimization. The final recommendation list is generated by calculating the scores defined in Eq. (2) on all items for each user. The recommendation scores on all items are sorted in descending order to produce the top-$N$ list.
5. Experiments
We conduct several empirical studies to demonstrate the usefulness of the SPMRec and its components. We also analyze the improvements for a better understanding of the underlying mechanism. Specifically, we answer the following six research questions (RQs):
  • RQ1: Does SPMRec provide more satisfactory recommendations than state-of-the-art baselines?
  • RQ2: How does our item side spectrum smoothing $\beta$ affect the diversity and recommendation performance?
  • RQ3: Does SPMRec outperform existing degeneration solutions?
  • RQ4: How does each component of SPMRec contribute to the performance?
  • RQ5: Where do improvements of SPMRec come from?
Table 1. Datasets Statistics.
Dataset #users #items #interactions density
avg.
interactions
per user
Home 66,519 28,237 551,682 0.03% 8.3
Beauty 22,363 12,101 198,502 0.05% 8.3
Toys 19,412 11,924 167,597 0.07% 8.6
Tools 16,638 10,217 134,476 0.08% 8.1
Office 4,905 2,420 53,258 0.44% 10.8
Total 129K 74K 1.10M - -
5.1. Datasets
We present the statistics of the dataset in Table 1. We evaluate all models in four benchmarks public Amazon review datasets [^1] http://deepyeti.ucsd.edu/jianmo/amazon/index.html. We choose Beauty, Toys and Games (Toys), Tools and Home (Tools), and Office Products (Office) categories in our experiments as these four categories are widely adopted benchmark categories (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022). We treat the existence of user-item reviews as positive user-item interactions. For each user, we sort the interacted items chronologically to form the user interaction sequence. The last interaction of the user sequence is used for testing and the second to last one for validation. We adopt the standard 5-core pre-processing setting on users (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022) to filter out users with less than five interactions. The detailed dataset statistics are presented in Table 1.
5.2. Baselines and Hyper-parameter Grid Search
We compare the proposed SPMRec with state-of-the-art sequential recommendation methods. We only include the static method BPRMF (Rendle et al., 2012) due to the page limitation. For sequential methods, we compare Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018), DT4SR (Fan et al., 2021), BERT4Rec (Sun et al., 2019), FMLP-Rec (Zhou et al., 2022), and STOSA (Fan et al., 2022). DuoRec (Qiu et al., 2022) addresses the degeneration issue in contrastive learning for SR. Note that we use only one training negative sample (i.e., $N^{-}_{u}=1$) for models with the Cross-Entropy loss for fair comparisons (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021).
5.3. All Items Ranking and Diversity Evaluation
We generate the top-N recommendation list for each user based on the dot-product between sequence output embedding and the entire item set in descending order. We rank all items for all models so that no item sampling bias is introduced in evaluation for fair comparisons (Krichene and Rendle, 2020). The evaluation includes classical top-N ranking metrics, Recall@N and NDCG@N. We present the averaged results over all test users. The test results are reported based on the best validation results. We report metrics in multiple Ns, including $N=\{5,10,40\}$, which are widely adopted in most existing SR methods (Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2022).
We evaluate the diversity of the recommendation list for all test users in two widely adopted metrics, including the Intra-list diversity (Zhang and Hurley, 2008; Ziegler et al., 2005) and the Coverage@100 of item categories for each user’s top-N recommendation list. We focus on the top-10 ranking list for intra-list diversity calculation and top-100 for the coverage metric. The intra-list diversity is calculated as follows:
(15) $Diversity(u)=1-\frac{\sum_{i\in\mathcal{I}_{u}^{N}}\sum_{j\in\mathcal{I}_{u}^{N}}\text{cos}(\mathbf{M}_{i},\mathbf{M}_{j})}{|\mathcal{I}_{u}^{N}|*|\mathcal{I}_{u}^{N}|},$
where $\mathcal{I}_{u}^{N}$ denotes the top-$N$ recommendation list for the user $u$, $\text{cos}(\cdot,\cdot)$ represents the cosine similarity, $\mathbf{M}$ is the item embedding table defined in Section 2.2, and the reported diversity is the average diversity over all users. The larger value of diversity indicates higher intra-list diversity for the recommendation. Note that, we also found similar improvements in other larger $N$s, such as $N=\{20,40\}$.
Table 2. Overall Performance Comparison Table. The best results are bold and the best baseline results are underlined, respectively. ‘Imp. vs. SASRec’ represents the relative improvement against SASRec without spectrum smoothing. ‘Improve.’ indicates the relative improvement against the best baseline performance.
Dataset Metric BPRMF Caser SASRec DT4SR BERT4Rec FMLP-Rec STOSA DuoRec SPMRec Imp. vs. SASRec Improv.
[…]
Cov@100 72.644 OOM 74.799 76.041 75.663 79.518 78.881 75.776 76.676 +2.51% -
[…]
Cov@100 44.881 45.147 45.066 46.127 45.668 47.929 47.333 47.199 48.050 +6.62% +0.03%
[…]
Cov@100 60.781 63.473 62.871 67.134 66.743 68.545 67.441 67.851 68.212 +8.50% -
Toys Recall@5 0.0301 0.0240 0.0551 0.0550 $0$ul0.0625 0.0577 0.0580 0.0631 +14.52% +0.96%
[…]
Cov@100 42.981 44.217 43.444 45.662 44.891 49.507 46.189 46.771 47.572 +9.50% -
[…]
Cov@100 49.887 50.243 51.299 51.341 51.001 51.335 51.661 51.514 52.733 +2.87% +2.08%
5.4. Overall Comparisons (RQ1)
We present the overall performance comparison with relevant state-of-the-art baselines in Table 2. From Table 2, we can easily conclude the superiority of SPMRec in providing sequential recommendations in various datasets and achieving not only better performance but also more diversity. We have several observations to conclude as follows:
  • SPMRec achieves the best performance and diversity, and outperforms SASRec with large improvement margins. In most performance metrics and all datasets, SPMRec significantly outperforms the second-best baseline with 0.52% to 12.33% in ranking performance and 4.71% to 14.41% in diversity. The improvements demonstrate the effectiveness of SPMRec in the sequential recommendation, while the recommendation diversity is also noticeably improved. Furthermore, as SPMRec is built on top of the original implementation of SASRec (Kang and McAuley, 2018), we observe 6.92% to 22.53% relative improvements over SASRec. We argue that the improvements originate from the singular spectrum smoothings on both sequence and item sides.
  • SPMRec can achieve better balance between recommendation quality and diversity. Comparing the diversity of SPMRec and baselines, we can observe that SPMRec consistently achieves better diversity results while having better performances, indicating the better capability of SPMRec in balancing the recommendation quality and diversity.
  • SPMRec achieves performance larger improvements in sparser datasets. Comparing SPMRec and SASRec, we can observe at least 17% improvements in both Beauty and Tools categories datasets, which have lower densities of 0.05% and 0.07%, respectively. The recommendation diversity is also significantly improved, especially for tools with a 14.43% improvement. However, in Toys and Office datasets with higher densities of 0.08% and 0.44%, the improvements are at most 14%. This observation demonstrates that the proposed singular spectrum smoothing SPMRec benefits sparser datasets, especially for datasets with a large portion of short sequences.
  • The consideration of Frobenius norm benefits the recommendation and diversity. STOSA adopts the sum of two L2 distances on mean and covariance embeddings and DuoRec utilizes the uniformity loss based on the exponential L2 distance. We can observe that STOSA and DuoRec both outperform the dot-product-based SASRec framework in recommendation quality and diversity. This observation further demonstrates the need to consider the Frobenius norm in embedding learning. Moreover, SPMRec performs better than STOSA and DuoRec, demonstrating the necessity of considering the regularization from the singular spectrum perspective.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 4. Relationship between item spectrum smoothing weight $\beta$ and diversity based on embeddings.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 5. Relationship between item spectrum smoothing weight $\beta$ and diversity based on coverage@100.
5.5. Diversity and Item Smoothing $\mathcal{L}_{\text{item}}$ (RQ2)
Item Spectrum Smoothing controlled by $\beta$ is closely related to the recommendation diversity. We demonstrate a strong connection between our item spectrum smoothing and the recommendation diversity by showing the relationship of $\beta$ and the diversity, as shown in Fig. (4). In Fig. (4), the x-axis denotes the value of $\beta$. As the $\beta$ becomes larger, we can observe that the diversity consistently increases, in all four datasets, demonstrating the strong connection between the recommendation diversity and our proposed item spectrum smoothing component.
Moreover, the recommendation performance drops when we demand too high recommendation diversity. Another observation is that the NDCG@10 drops significantly when we set $\beta$ to a too-large value, which also indicates a higher diversity. This observation demonstrates that a proper selection of $\beta$ can achieve satisfactory recommendation performance and diversity simultaneously. However, demanding higher diversity can hurt the recommendation performance.
Table 3. Performance Comparison Table for various methods to alleviate degeneration issues.
Dataset Metric SASRec SASRec+CosReg SASRec+Euclidean SPMRec
Beauty Recall@5 0.0416 0.0410 0.0326 0.0517
NDCG@5 0.0274 0.0250 0.0212 0.0345
Recall@10 0.0633 0.0609 0.0487 0.0745
NDCG@10 0.0343 0.0314 0.0264 0.0418
Tools Recall@5 0.0284 0.0290 0.0230 0.0350
NDCG@5 0.0194 0.0194 0.0145 0.0238
Recall@10 0.0427 0.0435 0.0359 0.0513
NDCG@10 0.0240 0.0240 0.0186 0.0290
Toys Recall@5 0.0551 0.0457 0.0537 0.0631
NDCG@5 0.0377 0.0302 0.0375 0.0431
Recall@10 0.0797 0.0658 0.0717 0.0852
NDCG@10 0.0465 0.0367 0.0432 0.0503
Office Recall@5 0.0656 0.0612 0.0561 0.0714
NDCG@5 0.0428 0.0400 0.0376 0.0489
Recall@10 0.0989 0.0919 0.0862 0.1036
NDCG@10 0.0534 0.0500 0.0472 0.0593
5.6. Degeneration Methods Comparison (RQ3)
We compare the proposed SPMRec with spectrum smoothing regularization with existing methods for alleviating the degeneration issue in word embeddings learning, including the cosine regularization (Gao et al., 2019) as follows:
(16) $\mathcal{L}_{\text{item}}=\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\frac{\mathbf{M}_{i}^{\top}\mathbf{M}_{j}}{||\mathbf{M}_{i}||\times||\mathbf{M}_{j}||},$
and the Euclidean method (Zhang et al., 2020) as follows:
(17) $\mathcal{L}_{\text{item}}$ $=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{M}_{i}-\mathbf{M}_{j}||_{2}.$
We apply the cosine and Euclidean regularization methods to both sides. We report the best grid searched results of all regularization methods in Table 3. Note that the uniformity property adopted by DuoRec and DirectAU (Qiu et al., 2022; Wang et al., 2022b) falls into the Euclidean methods category.
As shown in Table 3, neither cosine regularization or Euclidean regularization can consistently outperform the original SASRec without any regularization. Arguably, the potential reason is that neither of these two methods constrains the Frobenius norm of embeddings, which is the upper bound of the largest singular value, as discussed in Section 3.1. When cosine regularization and the Euclidean method are minimized, the naive solution is $||\mathbf{E}_{i}||_{2}\rightarrow\infty$ for all $i$ if we do not consider the supervised recommendation loss. Moreover, different from the word embedding task, sequences in the recommendation task are shorter on average, posing a greater challenge in addressing the degeneration in SR. However, our proposed singular spectrum smoothing method SPMRec minimizes the Frobenius norm while maximizing the nuclear norm, achieving the goals of smoothing the singular value curve and suppressing the largest singular value simultaneously.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 6. NDCG@5 on different sequential spectrum smoothing weight $\lambda$ values. Best Viewed in colors.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 7. NDCG@5 on different item spectrum smoothing weight $\beta$ values.
5.7. Ablation Study (RQ4)
We conduct the ablation study to show the effectiveness of two important components, including singular spectrum smoothing for sequence output and item embeddings. We investigate the smoothing weights $\lambda$ (for sequence output embeddings smoothing) and $\beta$ (for item embeddings smoothing) in Fig. (6) and Fig. (7), respectively. We include the SASRec (without any regularization) and DuoRec (with uniformity regularization). Note that the special case ($\lambda=0$ and $\beta=0$) is included, which is SASRec without any regularization. Additionally, the sequence outputs and item embeddings spectrum smoothing losses are in different scales ($10^{-1}$ for sequence embeddings and $10^{-4}$ for item embeddings). Hence, the scales of $\lambda$ and $\beta$ are different with the purpose of balancing the contributions of two regularizations.
5.7.1. Sequence Output Embedding Smoothing Weight $\lambda$
We present the performance values with different values of $\lambda$ in Fig. (6) for all four datasets. Fig. (6) shows that the proper selection of $\lambda$ improves the performance significantly. Moreover, most $\lambda$ values achieve comparative and better performances than SASRec. These observations show that the sequence output embedding spectrum smoothing is beneficial for the sequential recommendation.
5.7.2. Item Embedding Smoothing Weight $\beta$
The NDCG@5 values of different $\beta$ are shown in Fig. (7), which is controlling the spectrum smoothing component’s contribution. The analysis is conducted in all four datasets used. Overall, when $\beta$ becomes larger, i.e., having a larger contribution to the overall loss, the performance increases. It shows the effectiveness of item embeddings singular spectrum smoothing for performance improvements.
5.8. Improvements Analysis (RQ5)
5.8.1. Improvements w.r.t. Sequence Length
SPMRec is highly effective in short sequences recommendation. We group sequences (users) based on the number of training interactions (sequence length) of users. Within each group, we average the testing performances of sequences in the group. The comparisons on each group are shown in Fig. (8). We can observe that SPMRec achieves significant improvements in short sequences. Specifically, the improvements over the best STOSA range from 7.21% to 23.04%. The recommendation for short sequences is challenging due to limited interaction signals, which is known as the cold start problem. SPMRec regularizes the sequence output embeddings and prevents the degeneration so that short sequences can be distinguishable and are properly modeled. We can also observe that SPMRec has inferior performances in the longest sequence group, which is even worse than SASRec in the Toys dataset. However, the short sequences are in majority in both the used datasets and practical applications. With the proposed singular spectrum smoothing, SPMRec benefits short sequences with the sacrifices of long sequences.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 8. NDCG@5 on different sequence groups based on sequence length.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 9. NDCG@5 on different item groups based on item popularity.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 10. Normalized Singular Values Curve Comparison between SASRec and the proposed SPMRec in best grid searched settings.
5.9. Sequential User Degeneration Analysis (RQ5)
We visualize the singular values curve of sequence output embedding for our proposed SPMRec and SASRec, as shown in Fig. (10). We can observe that the proposed SPMRec has a smoother decaying curve than SASRec. Moreover, for SPMRec, the degrees of spectrum smoothness for all four datasets are 26.34 for Beauty, 20.76 for Tools, 20.25 for Toys, and 19.24 for Office. However, SASRec has lower degrees for spectrum smoothness as 19.54 for Beauty, 17.10 for Tools, 17.00 for Toys, and 13.67 for Office. SPMRec consistently has larger degrees of spectrum smoothness than SASRec, which demonstrates the effectiveness of SPMRec in preventing sequences from the degeneration for the sequential recommendation.
6. Related Work
6.1. Sequential Recommendation
Sequential Recommendation (SR) is a technique that encodes the temporal interactions of each user as a sequence and uses the encoded behaviors to infer their preferences. The main idea behind SR is to model the transitions between items within sequences. The earliest SR methods used Markovian approaches, such as FPMC (Rendle et al., 2010) and Fossil (He and McAuley, 2016), with the assumption that the item prediction depends only on the previous interacted items Another line of work utilizes the Recurrent Neural Network (RNN) as the base model for modeling sequential interactions. Several works adopt RNNs to address the SR problem, including GRU4Rec (Hidasi et al., 2015), HGN (Ma et al., 2019; Peng et al., 2021), HRNN (Quadrana et al., 2017), and RRN (Wu et al., 2017). Different variants of RNNs are proposed, such as the original GRU in GRU4Rec and hierarchical RNN in HGN and HRNN. The third line of methods to solve the SR problem is using the Convolution Neural Network (CNN), with the interpretation of a sequence of item embeddings as an image, including Caser (Tang and Wang, 2018), CosRec (Yan et al., 2019), and NextItNet (Yuan et al., 2019). The recent success of the self-attention module (Vaswani et al., 2017) inspires the adaptation of Transformer to the sequential recommendation task. The self-attention module captures the sequential high-order item transitions by modeling attention from all items in the sequence. SASRec (Kang and McAuley, 2018) is the earliest work of applying Transformer to SR. BERT4Rec (Sun et al., 2019) builds upon BERT (Devlin et al., 2019) with bidirectional Transformers for SR. FMLP-Rec (Zhou et al., 2022) proposes an all-MLP architecture in the Transformer for SR. Several variants of Transformers (Fan et al., 2021; Li et al., 2020; Wu et al., 2020; Liu et al., 2021) are further proposed.
6.2. Representation Degeneration
The representation degeneration problem describes the issue of learned embeddings being narrowly distributed in a small cone (Gao et al., 2019). The degeneration issue is firstly proposed in the word representations learning problem. As shown in (Gao et al., 2019), words from text data follow the long-tail distribution, which degenerates the representation learning process. Theoretically, it causes the narrow distribution of words in the representation latent space (Gao et al., 2019). Moreover, an empirical analysis (Yu et al., 2022) shows that the learning of tail words degenerates the representation optimization of frequent words. An important analysis in (Wang et al., 2020) shows that degeneration is closely related to the fast decaying singular values distribution. In other words, the largest singular value is significantly larger than all other singular values.
Distinct from the natural language processing (NLP) area, SR still has several unique characteristics. For example, sequences are short (data sparsity in recommendation systems), while sentences/paragraphs in NLP are typically long. This distinct characteristic introduces additional challenges for SR, where the degeneration issue can also happen in the sequence output embeddings. DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec and DirectAU (Wang et al., 2022b) optimizes the uniformity property in the in-batch representations to enforce the representations to be uniformly distributed, where the uniformity is based on the Euclidean distance. This work connects the item embeddings degeneration with the degeneration of sequence representations. The intertwining of both sequence and item rank degeneration is rarely discussed before in SR. Moreover, some solutions from NLP to the degeneration problem only apply to the item side, including the CosReg (Gao et al., 2019) based on cosine similarities, Euclidean regularization based on learned word similarities (Zhang et al., 2020), spectrum control (Wang et al., 2020; Yan et al., 2022), and gradient gating (Yu et al., 2022).
As discussed in (de Souza Pereira Moreira et al., 2021), natural language processing (NLP) and sequential recommendation (SR) share multiple similarities, including both learning discrete tokens learning (words in NLP and items in SR) and long-tail distribution of frequency for both words and items. With this inspiration, DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec optimizes the uniformity property in the in-batch augmented representations to alleviate the degeneration issue in contrastive learning, where the uniformity is based on the Euclidean distance.
7. Conclusion
We theoretically and empirically identify the relationship between the representation degeneration and recommendation diversity. Motivated by this connection, we propose a novel metric to measure the degree of degeneration as the area under the singular value curve (AUSC) to constrain the sequence and item embeddings learning. We further propose a singular spectrum smoothing regularization as a surrogate of AUSC to alleviate the degeneration and improve the diversity simultaneously. We empirically demonstrate the effectiveness of the proposed singular spectrum smoothing regularization in recommendation performances, especially for short sequences and popular items. We also empirically demonstrate a strong connection between the proposed singular spectrum smoothing regularization and the recommendation diversity.
Appendix A Baselines and Hyper-parameters Grid Search
We implement SPMRec with Pytorch based on the S3-Rec (Zhou et al., 2020) code base. We grid search all parameters and report the test performance based on the best validation results. We search the representation dimension in $\{64,128\}$, maximum sequence length from $\{50,100\}$, learning rate in $\{10^{-3},10^{-4}\}$, the $L2$ regularization control weight from $\{0.0,10^{-1},10^{-2},10^{-3}\}$, dropout rate from $\{0.3,0.5,0.7\}$ for all methods. For sequential methods, we search number of layers from $\{1,2,3\}$, and number of heads in $\{1,2,4\}$. We adopt the early stopping strategy that model optimization stops when the validation MRR does not increase for 50 epochs. The followings are the model specific hyper-parameters search ranges of baselines:
  • BPR [^2] https://github.com/xiangwang1223/neural_graph_collaborative_filtering: BPR is the most classical collaborative filtering method for personalized ranking with implicit feedbacks. We search the learning rate in $\{10^{-3},10^{-4}\}$ and $L2$ regularization weight from $\{10^{-1},10^{-2},10^{-3}\}$.
  • Caser [^3] https://github.com/graytowne/caser_pytorch: A CNN-based sequential recommendation method that views the sequence embedding matrix as an image and applies convolution operators to it. We search the length $L$ from $\{5,10\}$, and $T$ from $\{1,3,5\}$.
  • SASRec [^4] https://github.com/RUCAIBox/CIKM2020-S3Rec: The state-of-the-art sequential method that depends on the Transformer architecture. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • BERT4Rec [^5] https://github.com/FeiSun/BERT4Rec: This method extends SASRec to model bidirectional item transitions with standard Cloze objective. We search the mask probability from the range of $\{0.1,0.2,0.3,0.5,0.7\}$.
  • DT4SR [^6] https://github.com/DyGRec/DT4SR: A metric learning-base sequential method that models items as distributions and proposes mean and covariance Transformers. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • STOSA [^7] https://github.com/zfan20/STOSA: A metric learning-base sequential method that models items as distributions and proposes a Wasserstein self-attention module. We search the dropout rate from $\{0.3,0.5,0.7\}$.
  • DuoRec: [^8] https://github.com/RuihongQiu/DuoRec This method introduces unsupervised Dropout and supervised semantic augmentations in self-supervised learning for sequential recommendation.
  • FMLP-Rec: [^9] https://github.com/Woeee/FMLP-Rec Another most recent SR method adopting the full mlp layers architecture and filter methods on the self-attention module.